Consider the normal distribution of weights of the students in a college with mean 55kg and standard deviation 5kg . What is the probability that a student
i) weighs between 60.5 and 64 5. kgs ?
ii) weighs more than 65kgs ?
1
Expert's answer
2019-04-23T12:26:30-0400

i) P(60.5<X<64.5)=P((60.5-55)/5<Z<(64.5-55)/5)=P(1.1<Z<1.9)=

=P(Z<1.9)-P(Z<1.1)=0.9713-0.8643=0.1070.

ii) P(X>65)=P(Z>(65-55)/5)=P(Z>2)=1-P(Z<2)=1-0.9772=0.0228.